# Directed Acyclic Outerplanar Graphs Have Constant Stack Number

### v2

@@ 3.4 A Constant Upper Bound
Fix a base edge for $G$ and hence the corresponding construction tree. By Lemma 3.12 there is a directed $H$-partition $\mathcal{P}$ of $G$ satisfying properties (P1)(P4). In particular, by property (P3) the DAG $H$ is block-monotone. Thus by Corollary 3.6 there is an $h$-stack layout $\prec_{H}$ of $H$ with $h\leq 258$. Further, by property (P1), every part of $\mathcal{P}$ induces a transitive subgraph below some monotone vertex of $G$ and as such admits a $1$-stack layout by Lemma 3.14.
Now we seek to apply Lemma 3.10, for which we have to check its premises:
  • By property (P4), for every block $B$ of $H$ the directed $B$-partition $\mathcal{P}_{B}$ of $G[B]$ has cut cover number at most $w=4$. Thus, by Lemma 3.8, graph $G[B]$ admits an $H$-expanding stack layout using at most $s\leq 3wh+1=3\cdot 4\cdot 258+1=3097$ stacks.
  • By property (P2), in each part $P\in\mathcal{P}$ there are $p=2$ directed paths $Q_{1},Q_{2}$ in $G[v_{P}]$ such that all vertices stacked onto edges of $G[v_{P}]$ are stacked onto an edge of $Q_{1}$ or $Q_{2}$. As $G$ is outerplanar, we additionally get that at most $t=1$ vertex is stacked onto each of those edges.
Seeing all premises fulfilled, Lemma 3.10 yields that the stack number of $G$ is at most $\operatorname{sn}(G)\leq 4spt\leq 4\cdot 3097\cdot 2\cdot 1=24776$.
@@ 4 Directed Acyclic 2-Trees have Unbounded Stack Number
Consider well-interleaved triangles $T_{i}=(u_{i},v_{i},w_{i})$ for $i=1,\dots,k^{3}$. Without loss of generality we have $u_{1}\prec\dots\prec u_{k^{3}}\prec v_{1},\dots,v_{k^{3}}\prec w_{1},\dots,w_{k^{3}}$, that is, only the ordering within the $v$-vertices and within the $w$-vertices is unknown. Among the $v$-vertices, the Erdős-Szekeres theorem yields an increasing sequence of $k$ indices, i.e., a $k$-twist between $u$- and $v$-vertices with which we are done, or a decreasing sequence of $k^{2}$ indices with which we continue. From now on, we only consider these $k^{2}$ indices. Again, by Erdős-Szekeres, there either is an increasing sequence of length $k$ among the $k^{2}$ considered $w$-vertices yielding a $k$-twist with the corresponding $u$-vertices. Or there is a decreasing sequence of length $k$ giving a $k$-twist between the $v$- and $w$-vertices.
See 1.4
@@
[…]
Construction of $\boldsymbol{G}$. We define the desired $2$-tree $G$ via a sequence of $2$-trees $G_{0}\subset G_{1}\subset\cdots\subset G_{k}$ with $G_{k}=G$. For each $t=0,\ldots,k$ we shall have a matching $E_{t}\subset E(G_{t})$ such that in $G_{t}$ no vertex is stacked onto any edge in $E_{t}$, and $E_{0},\ldots,E_{k}$ are pairwise disjoint.
We start with $G_{0}$ being a single edge $ab$ oriented from $a$ to $b$, and $E_{0}=\{ab\}$. Having defined $G_{t}$ and $E_{t}$ for some $0\leq t<k$, we define $G_{t+1}$ and $E_{t+1}$ as follows: We use each edge $ab\in E_{t}$ (directed from $a$ to $b$) as the base edge for a particular $2$-tree that we denote by $T(ab)$ and that is constructed as follows: Let $N$ be a large enough integer. (We give a precise lower bound for $N$ below.)
@@
See Figure 13 for an illustration.
Figure 13: Construction of $T(ab)$ with the edge set $E(ab)$ (orange)
Observe that $T(ab)$ involves no transitive stackings, that at most two vertices are stacked onto each edge of $T(ab)$, and that in $T(ab)$ no vertex is stacked onto an edge in $E(ab)$. Further observe that in every vertex ordering of $T(ab)$ the vertices $b^{1},\ldots,b^{N}$ come (actually in that order) to the right of $a$ and $b$, while the vertices $a^{1},\ldots,a^{N}$ come (not necessarily in that order) to the left of $a$ and $b$. In particular, $a$ and $b$ are consecutive in every vertex ordering of $T(ab)$.
@@
For our proof below to work, we require that $N$ is an enormous (but constant) number in terms of $k$. We set $N=r_{1}k$, where $r_{1},\ldots,r_{k}$ is a sequence of integers satisfying
[…]
@@
[…]
Twist number of $\boldsymbol{G}$. Let $\prec$ be an arbitrary vertex ordering of $G$. We give an inductive proof of a slightly stronger statement than the existence of a $k$-twist, for which we need the following definition. For positive integers $r,t$ we call a matching a $t$-twist with an $r$-thick edge if it is obtained from a $t$-twist by replacing one edge by an $r$-rainbow. In particular, the $r$-rainbow is vertex-disjoint from the remaining $(t-1)$-twist and each rainbow-edge crosses each twist-edge. Our long-term goal is to find such a $t$-twist with a thick edge in $G_{t}$. However, we may also find a $k$-twist along the way, in which case we can stop immediately. Thus, throughout the proof we always assume that we do not find a $k$-twist. Under this assumption, we now give our stronger statement that we prove by induction: For every $t=1,\dots,k$, the subgraph $G_{t}$ of $G$ contains a $t$-twist with an $r_{t}$-thick edge, where the $r_{t}$-rainbow consists of edges in $E_{t}$. We start with a huge rainbow, which rapidly decreases while increasing the size of the twist by 1 in each step until we obtain a $k$-twist for $t=k$.
@@
Now, for $t\geq 1$, assume that we have a $t$-twist $T$ with an $r_{t}$-thick edge, where $R\subseteq E_{t}$ denotes the $r_{t}$-rainbow. We aim to find an entirely new rainbow $R^{\prime}\subseteq E_{t+1}$ of size $r_{t+1}$ and an edge $e^{\prime}$ crossing all edges of $R^{\prime}$, where all these edges start in the region spanned by the startpoints of $R$ and end in the region spanned by the endpoints of $R$. Together with the $t-1$ edges of $T-R$, this forms a $(t+1)$-twist with an $r_{t+1}$-thick edge, see Figure 14.
@@
For this, recall that in the construction of $G$, the 2-tree $T(ab)$ is added to each edge $ab\in R\subseteq E_{t}$. To find $R^{\prime}$ and $e^{\prime}$, we follow two steps: First, we analyze the edges $ab\in R$ and their right children. Here, we either find a large subset of edges having their children far to the right (Figure 15 left), which yields well-interleaved triangles and thus a $k$-twist by Lemma 4.1. Or we have the other extreme that many edges of $R$ have their children close to their right endpoint (Figure 15 right). In the second step, we find $R^{\prime}$ and $e^{\prime}$ in the 2-tree $T(ab)$ of such an edge $ab$.
@@
Figure 15: Left: Three consecutive edges of the rainbow $R$ with their right children (blue vertices) skipping sufficiently many vertices of $R$ such that three well-interleaved triangles are formed. Right: Two edges of the form $ab\in S^{*}\subseteq R$ with the right children of $T(ab)$ immediately following $b$.
@@
Figure 16: Five consecutive edges of $R$ with their right children, edges to the left parent omitted for readability. Edges to children skipping more than one vertex of $R$ are drawn dashed. Each block of two consecutive edges of $R$ contains an edge whose child skips at most one vertex. The chosen subrainbow $S_{1}$ is highlighted thick.
@@
For the first part, consider an edge $ab\in R$ and the first right child $b^{1}$ in $T(ab)$, and count the number of vertices of $R$ that are skipped, i.e., the number of endpoints of edges of $R$ between $b$ and $b^{1}$. If there are $k^{3}$ consecutive edges $ab$ in $R$ such that $b^{1}$ skips at least $k^{3}$ vertices of $R$, then we obtain $k^{3}$ well-interleaved triangles (Figure 15 left) and therefore a $k$-twist by Lemma 4.1. As we are done in this case, we assume the other case in which we have a set $S\subseteq R$ with the following two properties: First, $S$ consists of at least $r_{t}/k^{3}$ edges whose first right child skips less than $k^{3}$ vertices of $R$. And second, among each $k^{3}$ consecutive edges of $R$ at least one belongs to $S$. Observe that among the edges of $S$, there is a subset $S_{1}\subseteq S$ of $r_{t}/(2k^{3})$ edges $ab$ such that the right child $b^{1}$ does not skip any vertex of $S_{1}$ (Figure 16).
@@
Next, consider the second right children, i.e., the vertex $b^{2}$ for edges $ab\in S_{1}$. With the same argument as above, we have one of the two extremes in Figure 15: Either $k^{3}$ consecutive edges of $S_{2}$ have their second right children far to the right forming well-interleaved triangles. Or there is a subrainbow $S_{2}\subseteq S_{1}\subseteq S\subseteq R$ of size $r_{t}/(2k^{3})^{2}$ such that for every edge $ab\in S_{2}$, its second right child $b^{2}$ does not skip any vertex of $S_{2}$. Repeating the argument $s$ times, we obtain a subrainbow of $S_{s}\subseteq\dots\subseteq S_{2}\subseteq S_{1}\subseteq R$ such that each edge $ab$ has many right children $b^{1},b^{2},\dots,b^{s}$ not skipping any vertex of $S_{s}$. Note that the size of the subrainbow shrinks by a factor of $2k^{3}$ in each step. Since $r_{t}=k^{3}(2k^{3})^{s}$ for $s=1+r_{t+1}k$, after $s$ steps we are left with a set $S_{s}\subseteq R$ of $k^{3}$ edges of the form $ab$ such that $b^{1}\prec\dots\prec b^{s}$ immediately follow $b$, i.e., no other vertex of $S_{s}$ or the considered right children is between $b$ and $b^{s}$, which concludes the first part.
@@
The second part considers the left children of the 2-trees $T(ab)$ to find $R^{\prime}$ and $e^{\prime}$. Observe that if all $k^{3}$ edges of $S_{s}\subseteq R$ have a left child outside of the region spanned by $R$, then these children together with their parents form $k^{3}$ well-interleaved triangles. As this yields a $k$-twist by Lemma 4.1, we may assume that there is an edge $ab\in S_{s}$ with the children $a^{1},\dots,a^{s-1}$ below the outermost edge of $R$, where again $s=1+r_{t+1}k$. Among these children, we find the startpoints of the new rainbow $R^{\prime}$ using the Erdős-Szekeres theorem: Either we find a sequence of $k$ increasing indices, then the respective $a^{i}b^{i}$-edges form a $k$-twist and we are done. Or we find a sequence of $r_{t+1}$ decreasing indices, then the respective $a^{i}b^{i}$-edges form our desired $r_{t+1}$-rainbow $R^{\prime}$. Finally, we choose $e^{\prime}=ab^{s}$ as an edge that crosses all edges of $R^{\prime}$, see Figure 17. Combing the $(t-1)$-twist $T-R$ with the edge $e^{\prime}$ and the rainbow $R^{\prime}\subseteq E_{t+1}$, we obtain a $(t+1)$-twist with an $r_{t+1}$-thick edge.
@@ 5 Conclusion and Open Problems
We proved that outerplanar DAGs have bounded stack number (Theorem 1.3) and that monotone $2$-trees with at most two vertices stacked on every edge have unbounded stack number (Theorem 1.4). In both cases we solved long-standing open problems or conjectures. In doing so we got pretty close to pinpointing the boundary between bounded and unbounded stack number of directed acyclic $2$-trees, hence DAGs of treewidth $2$. However, several interesting questions worth to be considered remain open.
@@
Our family of $2$-trees constructed to prove Theorem 1.4 is not upward planar. This motivates the following open problem to further narrow the gap between bounded and unbounded stack number.
Open Problem 5.2.

### v3

@@ 3.4 A Constant Upper Bound
Fix a base edge for $G$ and hence the corresponding construction tree. By Lemma 3.12 there is a directed $H$-partition $\mathcal{P}$ of $G$ satisfying properties 14. In particular, by property 3 the DAG $H$ is block-monotone. Thus by Corollary 3.6 there is an $h$-stack layout $\prec_{H}$ of $H$ with $h\leq 258$. Further, by property 1, every part of $\mathcal{P}$ induces a transitive subgraph below some monotone vertex of $G$ and as such admits a $1$-stack layout by Lemma 3.14.
Now we seek to apply 3.9, for which we have to check its premises:
  • By property 4, for every block $B$ of $H$ the directed $B$-partition $\mathcal{P}_{B}$ of $G[B]$ has cut cover number at most $w=4$. Thus, by Lemma 3.8, graph $G[B]$ admits an $H$-expanding stack layout using at most $s\leq 3wh+1=3\cdot 4\cdot 258+1=3097$ stacks.
  • By property 2, in each part $P\in\mathcal{P}$ there are $p=2$ directed paths $Q_{1},Q_{2}$ in $G[P]$ such that all vertices stacked onto edges of $G[P]$ are stacked onto an edge of $Q_{1}$ or $Q_{2}$. As $G$ is outerplanar, we additionally get that at most $t=1$ vertex is stacked onto each of those edges.
Seeing all premises fulfilled, 3.9 yields that the stack number of $G$ is at most $\operatorname{sn}(G)\leq 4spt\leq 4\cdot 3097\cdot 2\cdot 1=24776$.
@@ 4 Directed Acyclic 2-Trees have Unbounded Stack Number
For the proof, we use the Erdős-Szekeres theorem which states that every sequence of length $p\cdot q$ consisting of pairwise distinct integers contains a monotonically increasing sequence of length $p$ or a monotonically decreasing sequence of length $q$.
@@
Consider well-interleaved triangles $T_{i}=(u_{i},v_{i},w_{i})$ for $i=1,\dots,k^{3}$. Without loss of generality we have $u_{1}\prec\dots\prec u_{k^{3}}\prec v_{1},\dots,v_{k^{3}}\prec w_{1},\dots,w_{k^{3}}$, that is, only the ordering within the $v$-vertices and within the $w$-vertices is unknown. Among the $v$-vertices, the Erdős-Szekeres theorem with $p=k$ and $q=k^{2}$ yields an increasing sequence of $k$ indices, i.e., a $k$-twist between $u$- and $v$-vertices with which we are done, or a decreasing sequence of $k^{2}$ indices with which we continue. From now on, we only consider these $k^{2}$ indices. Again, by Erdős-Szekeres, there either is an increasing sequence of length $k$ among the $k^{2}$ considered $w$-vertices yielding a $k$-twist with the corresponding $u$-vertices. Or there is a decreasing sequence of length $k$ giving a $k$-twist between the $v$- and $w$-vertices.
[…]
@@
{subproof}
[Construction of $\boldsymbol{G}$.] We define the desired $2$-tree $G$ via a sequence of $2$-trees $G_{0}\subset G_{1}\subset\cdots\subset G_{k}$ with $G_{k}=G$. For each $t=0,\ldots,k$ we shall have a matching $E_{t}\subset E(G_{t})$ such that in $G_{t}$ no vertex is stacked onto any edge in $E_{t}$, and $E_{0},\ldots,E_{k}$ are pairwise disjoint.
We start with $G_{0}$ being a single edge $ab$ oriented from $a$ to $b$, and $E_{0}=\{ab\}$. Having defined $G_{t}$ and $E_{t}$ for some $0\leq t<k$, we define $G_{t+1}$ and $E_{t+1}$ as follows: We use each edge $ab\in E_{t}$ (directed from $a$ to $b$) as the base edge for a particular $2$-tree that we denote by $T(ab)$ and that is constructed as follows: Let $N$ be a large enough integer (to be specified below).
@@
See 4.3 for an illustration. {sidefigure} [Uncaptioned image] Construction of $T(ab)$ with the edge set $E(ab)$ (orange) Observe that $T(ab)$ involves no transitive stackings, that at most two vertices are stacked onto each edge of $T(ab)$, and that in $T(ab)$ no vertex is stacked onto an edge in $E(ab)$. Further observe that in every vertex ordering of $T(ab)$ the vertices $b^{1},\ldots,b^{N}$ come (actually in that order) to the right of $a$ and $b$, while the vertices $a^{1},\ldots,a^{N}$ come (not necessarily in that order) to the left of $a$ and $b$. In particular, $a$ and $b$ are consecutive in every vertex ordering of $T(ab)$.
@@
For our proof below to work, we require that $N$ is an enormous (but constant) number in terms of $k$. We set $N=r_{1}k$, where $r_{1},\ldots,r_{k}$ is a sequence of integers defined recursively by
$r_{k}$ $=1\qquad\text{and}$
@@
{subproof}
[Twist number of $\boldsymbol{G}$.] Let $\prec$ be an arbitrary vertex ordering of $G$. We give an inductive proof of a slightly stronger statement than the existence of a $k$-twist, for which we need the following definition. For positive integers $r,t$ we call a matching a $t$-twist with an $r$-thick edge if it is obtained from a $t$-twist by replacing one edge by an $r$-rainbow. In particular, the $r$-rainbow is vertex-disjoint from the remaining (t1)(t-1)-twist and each rainbow-edge crosses each twist-edge. Our long-term goal is to find such a $t$-twist with a thick edge in $G_{t}$. However, we may also find a $k$-twist along the way, in which case we can stop immediately. Thus, throughout the proof we always assume that we do not find a $k$-twist. Under this assumption, we now give our stronger statement that we prove by induction: For every $t=1,\dots,k$, the subgraph $G_{t}$ of $G$ contains a $t$-twist with an $r_{t}$-thick edge, where the $r_{t}$-rainbow consists of edges in $E_{t}$. We start with a huge rainbow, which rapidly decreases while increasing the size of the twist by 1 in each step until we obtain a $k$-twist for $t=k$.
@@
Now, for $t\geq 1$, assume that we have a $t$-twist $T$ with an $r_{t}$-thick edge, where $R\subseteq E_{t}$ denotes the $r_{t}$-rainbow. We aim to find an entirely new rainbow $R^{\prime}\subseteq E_{t+1}$ of size $r_{t+1}$ and an edge $e^{\prime}$ crossing all edges of $R^{\prime}$, where all these edges start in the region spanned by the starting points of $R$ and end in the region spanned by the endpoints of $R$. Together with the t1t-1 edges of TRT-R, this forms a $(t+1)$-twist with an $r_{t+1}$-thick edge, see Figure 9.
@@
For this, recall that in the construction of $G$, the 2-tree $T(ab)$ is added to each edge $ab\in R\subseteq E_{t}$. To find $R^{\prime}$ and $e^{\prime}$, we follow two steps: First, we analyze the edges $ab\in R$ and their right children. Here, we either find a large subset of edges having their children far to the right (4.3 left), which yields well-interleaved triangles and thus a $k$-twist by Lemma 4.1. Or we have the other extreme that many edges of $R$ have their children close to their right endpoint (4.3 right). In the second step, we find $R^{\prime}$ and $e^{\prime}$ in the 2-tree $T(ab)$ of such an edge $ab$.
{cfigure}[Uncaptioned image][Uncaptioned image]
@@
Left: Three consecutive edges of the rainbow $R$ with their first right children (blue vertices) skipping sufficiently many vertices of $R$ such that three well-interleaved triangles are formed. (Note that for simplification of the proof, all children are required to skip at least $k^{3}$ vertices of $R$, although this is not necessary and not shown here for better readability, e.g., the child of the outermost edge.) Right: Two edges of the form $ab\in S_{s}\subseteq R$ with the right children of $T(ab)$ immediately following $b$.
{cfigure}[Uncaptioned image]
@@
Five consecutive edges of $R$ with their first right children, edges to the left parent are omitted for readability. Edges to children skipping more than one vertex of $R$ are drawn dashed. Each block of two consecutive edges of $R$ contains an edge whose child skips at most one vertex. The chosen subrainbow $S_{1}$ is highlighted thick.
@@
For the first part, consider an edge $ab\in R$ and the first right child $b^{1}$ in $T(ab)$, and count the number of vertices of $R$ that are skipped, i.e., the number of endpoints of edges of $R$ between $b$ and $b^{1}$. If there are $k^{3}$ (along the nesting order) consecutive edges $ab$ in $R$ such that each $b^{1}$ skips at least $k^{3}$ vertices of $R$, then we obtain $k^{3}$ well-interleaved triangles (4.3 left) and therefore a $k$-twist by Lemma 4.1. As we are done in this case, we assume the other case in which we have a set $S\subseteq R$ with the following two properties: First, $S$ consists of at least $r_{t}/k^{3}$ edges whose first right child skips less than $k^{3}$ vertices of $R$. And second, among each $k^{3}$ consecutive edges of $R$ at least one belongs to $S$. It follows that among the edges of $S$, there is a subset $S_{1}\subseteq S$ of $r_{t}/(2k^{3})$ edges $ab$ such that the right child $b^{1}$ does not skip any vertex of $S_{1}$; choose one in each block of $2k^{3}$ consecutive edges in $R$ (4.3).
@@
Next, consider the second right children, i.e., the vertex $b^{2}$ for edges $ab\in S_{1}$. With the same argument as above, we have one of the two extremes in 4.3: Either $k^{3}$ consecutive edges of $S_{2}$ have their second right children far to the right forming well-interleaved triangles. Or there is a subrainbow $S_{2}\subseteq S_{1}\subseteq S\subseteq R$ of size $r_{t}/(2k^{3})^{2}$ such that for every edge $ab\in S_{2}$, its second right child $b^{2}$ does not skip any vertex of $S_{2}$. Repeating the argument $s$ times, we obtain a subrainbow of $S_{s}\subseteq\dots\subseteq S_{2}\subseteq S_{1}\subseteq R$ such that each edge $ab$ has many right children $b^{1},b^{2},\dots,b^{s}$ not skipping any vertex of $S_{s}$. Note that the size of the subrainbow shrinks by a factor of $2k^{3}$ in each step. Since $r_{t}=k^{3}(2k^{3})^{s}$ for $s=1+r_{t+1}k$, after $s$ steps we are left with a set $S_{s}\subseteq R$ of $k^{3}$ edges of the form $ab$ such that $b^{1}\prec\dots\prec b^{s}$ immediately follow $b$, i.e., no other vertex of $S_{s}$ or the considered right children is between $b$ and $b^{s}$, which concludes the first part.
@@
The second part considers the left children of the 2-trees $T(ab)$ to find $R^{\prime}$ and $e^{\prime}$. Observe that if all $k^{3}$ edges of $S_{s}\subseteq R$ have a left child outside the region spanned by $R$, then these children together with their parents form $k^{3}$ well-interleaved triangles. As this yields a $k$-twist by Lemma 4.1, we may assume that there is an edge $ab\in S_{s}$ with the children $a^{1},\dots,a^{s-1}$ below the outermost edge of $R$, where again $s=1+r_{t+1}k$. Among these children, we find the starting points of the new rainbow $R^{\prime}$ using the Erdős-Szekeres theorem: Either we find a sequence of $k$ increasing indices, then the respective $a^{i}b^{i}$-edges form a $k$-twist and we are done. Or we find a sequence of $r_{t+1}$ decreasing indices, then the respective $a^{i}b^{i}$-edges form our desired $r_{t+1}$-rainbow $R^{\prime}$. Finally, we choose $e^{\prime}=ab^{s}$ as an edge that crosses all edges of $R^{\prime}$, see Figure 10. Combing the (t1)(t-1)-twist TRT-R with the edge $e^{\prime}$ and the rainbow $R^{\prime}\subseteq E_{t+1}$, and recalling that the 2-tree $T(ab)$ including $R^{\prime}$ and $e^{\prime}$ is below the outermost edge of $R$, we obtain a $(t+1)$-twist with an $r_{t+1}$-thick edge.
@@ 5 Conclusion and Open Problems
We proved that outerplanar DAGs have bounded stack number (Section 1) and that monotone $2$-trees with at most two vertices stacked on every edge have unbounded stack number (Section 1). In both cases we solved long-standing open problems or conjectures. In doing so we got pretty close to pinpointing the boundary between bounded and unbounded stack number of directed acyclic $2$-trees, hence DAGs of treewidth $2$. However, several interesting questions worth to be considered remain open.
@@
Our family of $2$-trees constructed to prove Section 1 is not upward planar. This motivates the following open problem to further narrow the gap between bounded and unbounded stack number.
{open}